OpenStudy (anonymous):

i am in 7th grade and i need help with algebra? my math comprehension is not to good!!! can someone please show me how to solve 5y-6=8y+6 ytankyou

OpenStudy (mysesshou):

5y-6=8y+6 First, let's add six to each side: 5y-6=8y+6 +6 +6 this gives 5y=8y+6+6 simplified to 5y=8y+12 subtract 8y both sides 5y=8y+12 -8y -8y 5y-8y=12 -3y=12 divide both sides by -3 \[\frac{-3y}{-3}=\frac{12}{-3}\] \[y=\frac{12}{-3} = -4\] Here you go, hope this helps some :)

OpenStudy (mysesshou):

So, in conclusion, y = -4

OpenStudy (anonymous):

thank you

OpenStudy (mysesshou):

welcome ;)

OpenStudy (anonymous):

would the answer for 4a-10=3a-1 be a=9

OpenStudy (mysesshou):

yes, I posted that one in the other post :) http://openstudy.com/study#/updates/4f6b884ae4b014cf77c84c0f